OpenStudy (anonymous):

8 , 10 , 12 , 14, 16 , . . . express using explicit formula

OpenStudy (anonymous):

It is 2n + 6 where n is the term number for first term, n = 1 so 2(1) + 6 = 8 for second term, n = 2 so 2(2) + 6 = 10 for third term, n = 3 so 2(3) + 6 = 12 and so on